X=t-4
Y=t/2
Write as a single linear equation

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

X=t-4
Y=t/2
Write as a single linear equation
---------
t = x + 4 = 2y
x - 2y = -4

Answer by stanbon(75887)

X=t-4
Y=t/2
Write as a single linear equation
----------
t = x+4
t = 2y
------
2y = x+4
-------
y = (1/2)x + 2
-----------
Cheers,
Stan H.
-----------